# Early Classification of Time Series: Taxonomy of Methods and Extensive Benchmark

### v2

@@ 4.3 Experiments with unbalanced misclassification and non-linear delay costs
There are however applications where these assumptions do not hold, for instance predictive maintenance or hospital emergency services, are characterized by (i) imbalanced misclassification costs (e.g. it is more costly to have to repair a machine than to carry out a maintenance operation that turns out not being necessary) and by (ii) non linear delay cost (e.g. usually, the later the surgical operation is decided, the costlier it is to organize it and the larger the risk for the patient). In the following, we call all applications presenting these characteristics “anomaly detection” applications.
@@ 4.3.1 Cost definition for anomaly detection
In order to study the behavior of the various algorithms on scenarios corresponding to anomaly detection, we set the unbalanced misclassification cost matrix such that a false negative (i.e. missing an anomaly) was 100 times costlier than a false positive (i.e. wrongly predicting an anomaly) (see Figure 5(b)). For this last situation, the cost was arbitrarily set to 1. The delay cost is defined as an exponential function of time. In order to have a cost commensurable with the misclassification one, we decided that waiting for the entire time series to be seen, at $T$, would cost 100 (see Figure 5(a)), starting at 1 for $t=0$ and reaching 100 when $t=T$.
@@ 4.3.2 Results and analysis
In this part, as a new cost setting is explored, there is no need to produce comparable results from previous works. Thus, we choose to use the new non z-normalized datasets collection (orange cylinder in Figure 4). In order for the imbalanced misclassification cost to make sense, those datasets have been altered so that the minority class represents 20% of all labels. As explained in Section 4.1, some extrinsic regression datasets are turn into classification ones. In these cases, the threshold value has been set to the second decile of the regression target. For the original classification datasets, the minority class has been sub-sampled when necessary.
@@
Results from the Wilcoxon-Holm Ranked test (both regarding the average rank and the value for AvgCost) (see Figure 6(b)) and from the AvgCost plot (see Figure 6(a)) with varying values of $\alpha$ (in Equation 16) show that now the best method overall is Economy which is both cost-informed at training and testing time, beside being anticipating-based. However, Stopping-Rule is a very strong contender while being cost-informed at training time but not at testing time and confidence-based. There is a reason for it. When Stopping Rule equals or overpasses Economy, this is for high values of $\alpha$ when the delay cost loses its importance, therefore leaving the misclassification cost to reign and confidence-based methods to be good.
It may come as a surprise that Calimera lags behind Economy for $\alpha\in[0,0.4]$, despite being similarly based on the estimation of future cost expectation. One reason for this is that the cost expectation is achieved by considering only the predicted class. This poor estimate of the cost expectancy becomes critical when the delay cost is important.
Similarly, Proba Threshold is surprisingly good in this scenario, even if it is no longer in the top tier. Looking solely at prediction confidence, we might expect it to be blind to the rapid increase in delay cost in the anomaly detection scenario. However, it is noticeable that the cost of delay only increases sharply after around 60% of the complete time series has been observed, which is generally sufficient to exceed the confidence threshold. Hence, Proba Threshold does not suffer from high delay costs that are to come, and exhibits a good performance here.
Figure 7 plots the Pareto front considering two axes based on decision costs. The horizontal axis corresponds to the average delay cost incurred for each example, normalized by the worst delay cost paid at $t=T$. It is better to be on the left of the $x$-axis. The vertical axis corresponds to 1 minus the misclassification cost incurred for each example, normalized by the worst prediction cost. It is better to be high on the $y$-axis.
We observe that the Pareto front is composed almost exclusively of points corresponding to the Economy method. This is consistent with the evaluation based on the AvgCost metric. This figure highlights the fact that the design of approaches capable of handling arbitrarily parameterized decision costs requires a cost-informed application framework.
@@ 4.4 Other experiments: ablation and substitution studies
In this section, complementary experiments, namely ablation studies as well as sanity checks are briefly discussed. For the sake of brevity, the figures supporting the analysis are reported in Appendix B.
@@
? (?) assert that calibration of the classifiers is paramount for the performance of ECTS algorithms. In order to test this claim, we have repeated the experiments removing the calibration step. The examples used for calibration have been removed as well during training, so that all else remains the same as before.
The results of Figure 11 show that indeed Calimera suffers greatly if no calibration is done. Indeed, this approach relies on estimating the expectation of future costs via a regression problem, and a miscalibration may have a negative impact on the built targets. For its part, Proba Threshold suffers somewhat mildly. This is no surprise as they rely on a single threshold on the confidence of the prediction for all time steps.
@@
All methods have been compared using the same classifier: MiniROCKET so that only the decision components differ. However, the choice of the base classifiers could induce a bias favoring or hampering some methods. In order to clarify this, we have repeated the experiments replacing MiniROCKET with two base classifiers: WEASEL 2.0 (?), and the XGBoost classifier (?) using features produced by tsfresh (?). Both of these classifiers have already been tested within the ECTS literature by ? (?, ?) and ? (?) respectively. Figure 12 and 13 in Appendix B report the results respectively with these two classification methods. One can observe that the results are not significantly altered with the same overall ordering of the methods when varying the value of $\alpha$. Furthermore, our results on AvgCost show that performances tends to be better for all methods using MiniROCKET. It is thus to be preferred given its simplicity and good performances.
@@
Considering the newly proposed ensemble of datasets, we were not able to identify any problems of information leakage over time. This inconclusive result simply indicates that the variance of the time series measurements is not informative for theses datasets, which still could be the case considering past published results. For further details, please refer to Section B.5 of Appendix B.
@@ 5 Conclusion
The first contribution of this research work is the coding of all the methods tested and making the codes available in a repository open to everyone. In this way, the experiments reported can be duplicated and further studies carried out. Furthermore, the deposited datasets and the experimental framework provide a ground for fair comparisons bewteen competing methods. We claim that the AvgCost is the appropriate measure by which to evaluate the performance of the methods. This is indeed what will be “paid” at the end of the day by a practitioner using a method. We have accordingly characterized a number of methods from the literature.
Our extensive experiments have shown that:
  • It is worthwhile to resort to dedicated ECTS methods, and more so in scenarios like anomaly detection with asymmetrical misclassification cost and exponential delay cost.
  • However, it is noticeable that Proba Threshold, a baseline method, is surprisingly good overall in the standard setting with symmetrical misclassification cost and linear delay cost, exhibiting comparable performance to confidence-based myopic methods such as Stopping-Rule and anticipation-based cost-informed ones such as Calimera and Economy.
  • Calibration of the classifiers has a large impact on some methods such as Calimera in particular, less so on other methods like Proba Threshold and Ecdire.
In this paper, we have proposed a reading guide to highlight the main characteristics of ECTS methods, namely (i) the importance of the two components: decision and prediction which are distinct in the “separable” architecture and not in the “end-to-end” one, (ii) the distinction between anticipation-based and myopic methods, and (iii) between cost-informed and cost-uninformed techniques. On the basis of these dimensions, it becomes easy to imagine new methods that combine them in original ways, which could lead to new properties and better performance for solving the problem of early classification of time series, which, present in many applications has potentially great impacts.
@@
To go a step further, future work could be carried out to study the literature’s approaches applied in as yet unexplored cost settings. For example, in many applications, the delay cost depends on the true class and the predicted one, and thus a single cost function integrating misclassification and delay costs should then be used. The use of this general cost form requires the adaptation of some state-of-the-art methods and has not yet been studied.
In addition, in real ECTS applications, it is up to the business expert to define the costs, which is not an easy task in practice. Future work could study the impact of: (i) noisy costs, (ii) and cost drift between training and testing stages. This would make it possible to identify the most resilient literature approaches.
@@
A Data description
A.1 UCR Time Series Classification datasets

### v6

@@ 4.3 Experiments with unbalanced misclassification and non-linear delay costs
There are, however, applications where these assumptions do not hold, for instance, predictive maintenance or hospital emergency services, are characterized by (i) unbalanced misclassification costs (e.g. it is more costly to have to repair a machine than to carry out a maintenance operation that turns out not being necessary) and by (ii) non-linear delay costs (e.g. usually, the later the surgical operation is decided, the costlier it is to organize it and the larger the risk for the patient). In the following, we call all applications presenting these characteristics “anomaly detection” applications.
@@ 4.3.1 Cost definition for anomaly detection
In order to study the behavior of the various algorithms on scenarios corresponding to anomaly detection, we set the unbalanced misclassification cost matrix such that a false negative (i.e. missing an anomaly) was 100 times costlier than a false positive (i.e. wrongly predicting an anomaly) (see Figure 5(b)). For this last situation, the delay cost was arbitrarily set to 1. The delay cost is defined as an exponential function of time. In order to have a delay cost commensurable with the misclassification one, we decided that waiting for the entire time series to be seen, at $T$, would cost $100\times\alpha$ (see Figure 5(a)), starting at $(1-\alpha)$ for $t=0$ and reaching $100\times\alpha$ when $t=T$.
@@ 4.3.2 Results and analysis
In this part, as a new cost setting is explored, there is no need to produce comparable results from previous works. Thus, we choose to use the new non z-normalized datasets collection (orange cylinder in Figure 2). In order for the imbalanced misclassification cost to make sense, the datasets have been altered so that the minority class represents 20% of all labels. As explained in Section 4.1, some extrinsic regression datasets are turned into classification ones. In these cases, the threshold value has been set to the second decile of the regression target. For the original classification datasets, the minority class has been sub-sampled when necessary.
@@
Results from the Wilcoxon-Holm Ranked test (both regarding the average rank and the value for AvgCost) (see Figure 6(b)) and from the AvgCost plot (see Figure 6(a)) with varying values of $\alpha$ (in Equation 9) show that now, the best method overall is Economy which is cost-informed at testing time, in addition to being anticipation-based. However, Stopping-Rule is a very strong contender while being cost-informed but not at testing time and confidence-based. There is a reason for it. When Stopping Rule equals or overpasses Economy, this applies to high values of $\alpha$ when the delay cost loses its importance, therefore leaving the misclassification cost to reign and confidence-based methods to be effective.
It may come as a surprise that Calimera lags behind Economy for $\alpha\in[0,0.4]$, despite being similarly based on the estimation of future cost expectations. One reason for this is that the cost expectation is achieved by considering only the predicted class. This poor estimate of the cost expectancy becomes critical when the delay cost is important.
Similarly, Proba Threshold is surprisingly good in this scenario, even if it is no longer in the top tier. Looking solely at prediction confidence, we might expect it to be blind to the rapid increase in delay cost in the anomaly detection scenario. However, it is noticeable that the cost of delay only increases sharply after around 60% of the complete time series has been observed, which is generally sufficient to exceed the confidence threshold. Hence, Proba Threshold does not suffer from high delay costs that are to come, and exhibits good performance here.
Figure 7 plots the Pareto front, considering two axes based on decision costs. The horizontal axis corresponds to the average delay cost incurred for each example, normalized by the worst delay cost paid at $t=T$. It is better to be on the left of the $x$-axis. The vertical axis corresponds to one minus the misclassification cost incurred for each example, normalized by the worst prediction cost. It is better to be high on the $y$-axis.
We observe that the Pareto front is composed almost exclusively of points corresponding to the Economy and Calimera methods. This is consistent with the evaluation based on the AvgCost metric. This figure highlights the fact that the design of approaches capable of handling arbitrarily parameterized decision costs requires a cost-informed application framework.
@@ 4.4 Other experiments: ablation and substitution studies
In this section, complementary experiments, namely ablation studies as well as sanity checks are briefly discussed. For the sake of brevity, the figures supporting the analysis are reported in Appendix LABEL:app_results.
@@
Bilski & Jastrzebska (2023) assert that calibration of the classifiers is paramount for the performance of ECTS algorithms. In order to test this claim, we have repeated the experiment, removing the calibration step. The examples used for calibration have also been removed during training, so that all else remains the same as before.
The results of Figure LABEL:fig:calib_heatbar in Appendix LABEL:calib show that indeed Calimera suffers greatly if no calibration is done. Indeed, this approach relies on estimating the expectation of future costs via a regression problem, and miscalibration may have a negative impact on the built targets. For its part, Proba Threshold suffers somewhat mildly. This is no surprise, as they rely on a single threshold on the confidence of the prediction for all time steps.
Impact of removing non-myopia
Experiments have shown that anticipation-based approaches tend to outperform the myopic competitors; we further validate this by conducting an ablation study over the non-myopia of both Economy and Calimera.
Figure LABEL:fig:non_myopia in Appendix LABEL:non_myopic_ablation clearly shows that, for weak temporal costs, the anticipation-based property is critical, whereas myopic counterparts significantly fall behind original methods. Intuitively, for higher temporal costs, the anticipation-based property is less important as decisions tend to be taken earlier, and it is less necessary for the methods to anticipate the future.
@@
All methods have been compared using the same classifier: MiniROCKET so that only the decision components differ. However, the choice of the base classifier could induce a bias favoring or hampering some methods. In order to clarify this, we have repeated the experiments replacing MiniROCKET with two base classifiers: WEASEL 2.0 (Schäfer & Leser, 2023), and the XGBoost classifier (Chen et al., 2015) using features produced by tsfresh (Christ et al., 2018). Both of these classifiers have already been tested within the ECTS literature by Schäfer & Leser (2020); Lv et al. (2019) and Achenchabe et al. (2021) respectively. Figure LABEL:fig:weasel and LABEL:fig:xgb in Appendix LABEL:change_clf report the results respectively with these two classification methods. One can observe that the results are not significantly altered with the same overall ordering of the methods when varying the value of $\alpha$. Furthermore, our results on AvgCost show that performance tends to be better for all methods using MiniROCKET (see Table LABEL:tab:clf_perfs in Appendix LABEL:change_clf). It is thus to be preferred given its simplicity and good performance.
@@
Considering the newly proposed ensemble of datasets, we were not able to identify any problems of information leakage over time. This inconclusive result simply indicates that the variance of the time series measurements is not informative for these datasets, which still could be the case considering past published results. For further details, please refer to Appendix LABEL:z_norm_app.
@@ 5 Conclusion
In this paper, we have proposed a taxonomy that allows to underline the main families of separable approaches for the ECTS problem, pointing out the essential components and the questions that have to be tackled when designing a new method. We have thus enlightened (i) the importance of the two components: decision and prediction, (ii) the distinction between anticipation-based and myopic methods, and (iii) between cost-informed and cost-uninformed techniques.
We have defined a methodology for evaluating and comparing ECTS methods. In addition, and we hope this will prove useful to the community, we have built an open source library that includes systematic implementation of the methods tested, the proposed evaluation protocol, as well as a collection of 34 datasets designed to enable informative testing of the methods.
We have also underlined the importance of considering a variety of cost settings in the evaluation of ECTS methods so as to reflect what can happen in real-life applications.
The in-depth experiments carried out shed light on design choices. Firstly, are anticipation-based methods better than myopic ones? We have shown that four of the eleven methods tested perform significantly better. However, these methods do not share the same characteristics with regard to this question. Secondly, to the question: do cost-informed methods outperform cost-uninformed methods? The Pareto front indicates a clear superiority of the former. Finally, when we test the robustness of methods to variations in cost functions, anticipation-based methods prove superior and, among them, cost-based methods fare even better.
Our experiments have also shown that calibration of the classifiers has a large impact on some methods (e.g. Calimera) in particular, less so on other methods (e.g. Proba Threshold and Ecdire). As regards to the z-normalization of time series which would be unwisely used in evaluation studies, our results (see Appendix LABEL:z_norm_app) show that the impact on the performance is limited, thus suggesting that the existing trigger functions do not or only slightly benefit from this information leakage.
@@
Future work could be carried out to study the literature’s approaches applied in as yet unexplored cost settings. For example, in many applications, the delay cost depends on the true class and the predicted one, and thus a single cost function integrating misclassification and delay costs should then be used. This general cost form requires the adaptation of some state-of-the-art methods and has not yet been studied.
In addition, in real ECTS applications, it is up to the business expert to define the costs, which is not an easy task in practice. Among the challenges, applications where the costs actually paid are not deterministic are of key interest (e.g. a manufacturing defect on an engine part does not necessarily lead to a failure, but it does increase the probability of paying a higher cost). Thus, future work could study the impact of stochastic cost functions. Another interesting case is applications where the costs paid are slightly different, or changed, from those defined by the business experts for the training phase (e.g. a change in the price of raw materials). Those kinds of cost drift between training and testing stages could also be further studied.
@@
Beyond separable approaches, a more extensive review of end-to-end ECTS methods could be interesting to complete the one presented in this paper. In particular, an experimental protocol should be designed to place end-to-end and separable methods on similar ground and to design new, efficient approaches.
Appendix A Benchmarking: related studies
Table 3: Table of published methods for the ECTS problem with their benchmarking protocol.
References No. of datasets No. of methods Stats test Eval. metric(s) Cost balance Code
Reject (Hatami & Chira, 2013) 1 1 Acc./Earl. fixed
iHMM (Antonucci et al., 2015) 1 2 $\ell_{0/1}$ fixed
ECDIRE (Mori et al., 2017b) 45 5 Acc./Earl. fixed
Stopping Rule (Mori et al., 2017a) 45 5 Acc./Earl. vary
ECEC (Lv et al., 2019) 45+8 6 Acc./Earl. vary
TEASER (Schäfer & Leser, 2020) 45 5 Acc./Earl./HM vary
SOCN (Lv et al., 2023) 45 10 Acc./Earl./HM fixed
ECTS (Xing et al., 2012) 23 3 Acc./Earl. fixed
RelClass (Parrish et al., 2013) 15 2 Acc./Earl. vary
2step/NoCluster (Tavenard & Malinowski, 2016) 76 3 AvgCost/Acc. vary
ECONOMY-$\gamma$-max (Zafar et al., 2021) 45/34 2 AvgCost vary
CALIMERA (Bilski & Jastrzebska, 2023) 45/34 8 AvgCost/Acc./Earl./HM fixed
FIRMBOUND (Ebihara et al., 2025) 7 5 AAPR/SAT vary
EDSC (Xing et al., 2011) 7 4 Acc./Earl. fixed
EARLIEST (Hartvigsen et al., 2019) 4 3 Acc./Earl. vary
DDQN (Martinez et al., 2020) 1 2 Acc./Earl. vary
DETSCNet (Chen et al., 2022) 12 5 HM fixed
Benefitter (Shekhar et al., 2023) 14 7 Acc./Earl. vary
CIS (Cao et al., 2023) 3 3 Acc./Earl. fixed
ELECTS (Rußwurm et al., 2023) 4 2 Acc./Earl. vary
EarlyStop-RL (Wang et al., 2024) 1 4 FPR/FNR/Earl./F1 fixed
Empirical Survey (Akasiadis et al., 2024) 3 5 Acc./Earl./HM/F1 fixed
Columns are defined as :
  • No. of datasets: the number of datasets used in the paper.
  • No. of methods: the number of methods tested in the paper (including the one proposed, if any).
  • Stats test: whether evaluation has been performed using statistical tests (Rainio et al., 2024).
  • Eval. metric(s): evaluation metrics used.
  • Cost balance: whether experiments have been conducted for different cost balances between $C_{m}$ and $C_{d}$.
  • Code: is the code of the method available.
Experimental methodology is heterogeneous in the literature so far, making comparisons between methods more difficult. When considering datasets used or evaluation metrics, a variety of different approaches can be found, as shown in Table 3.
Appendix B Experimental protocol
B.1 Optimization of the parameters of the methods
[…]
B.2 Hyperparameters
Table 4: Hyperparameters’ value. $\Delta$ defines grid’s size when performing grid-search over continuous valued intervals.
Hyperparameters
Method Fixed Optimized
Calimera kernel: "rbf"
[…]
Ects support = 0
[…]
Proba Threshold $\Delta=40$
Stopping Rule $\gamma_{1},\>\gamma_{2},\>\gamma_{3}\in[-1,1]$
$\Delta=10^{3}$
[…]
B.3 Calibration of the classifications
Like Bilski & Jastrzebska (2023), we add a calibration step when learning the classifiers, i.e. Platt’s scaling (Platt et al., 1999). Indeed, as we are dealing with collections of independently trained classifiers, the prediction scores may not remain consistent with one another over the time dimension. However, the trigger methods usually have their parameters set with the same values for all time steps. This is the case, for example with the Proba Threshold approach. In addition, some approaches such as Calimera and Economy-$\gamma$-Max exploit the estimated posterior probabilities $\{p(y|\textbf{x}_{t})\}_{y\in\mathcal{Y}}$ to estimate the future cost expectation. It is therefore highly desirable for all classifiers, at all times, to have their output calibrated and is necessary for a fair comparison.
B.4 Datasets selection
When a time series does not bring an increasing level of information over time about its class, classifiers are likely to obtain the same confusion matrix for all time steps, and the trigger function should then choose to make a prediction at the first time step to avoid the delay cost even if the prediction is very uncertain. Although this case may occur, it cannot form the basis for comparing the trigger functions. This is why we have been careful about the selection of data sets.
We have chosen data sets where the classification performance tends to increase over time, as measured with the same basis classifier for all methods (see Section 4.1.2). We check that the train AUC increases averaged over a time window of size 5, when using either the beginning of the series, or almost complete ones. Specifically, the beginning of the series is taken using $5\%,\ldots,25\%$, and the almost complete series are of the following lengths $75\%,\ldots,100\%$. When the difference of averaged AUC is strictly positive, the dataset is selected. (see Table LABEL:tab:orange_cylinder). In addition, all datasets exhibiting z-normalization have been removed. Thirty-four datasets passed these demanding criteria (see Appendix E).
For instance, Figure 8 (left) displays the time series of a synthetic data set where each of the three classes is characterized by a specific pattern over time. Class 0 can thus be recognized early on, while class 1 and class 2 can only be discriminated after approximately one third of the length of the time series. This is confirmed by the accuracy of the recognition of each class (right).
Selecting data sets appropriate for the comparison of ECTS methods is a demanding process that must check for all the desirable properties described above. We thus hope that the corpus thus built and made available on the open source library will be useful to the scientific community for future performance studies.
Figure 8: Examples of training time series of the SmoothSubspace dataset. Each of the class has a discriminative signal within one third of the serie (see Appendix D for a code snippet reproducing these results, based on the proposed library).
Appendix C End-to-end ECTS
C.1 State-of-the-art
Table 5: Table of published end-to-end methods for the ECTS problem.
References Architecture RL DL Cost-genericity
EDSC (Xing et al., 2011) Shapelet
EARLIEST (Hartvigsen et al., 2019) LSTM
DDQN (Martinez et al., 2020) MLP
DETSCNet (Chen et al., 2022) TCN
Benefitter (Shekhar et al., 2023) LSTM
CIS (Cao et al., 2023) LSTM
ELECTS (Rußwurm et al., 2023) LSTM
EarlyStop-RL (Wang et al., 2024) MLP
$\bullet$ A different class of methods relies on searching telltale representations of subsequences, such that if the incoming time sequence $\mathbf{x}_{t}$ matches one or more of these representations, then its class can be predicted. Typically, these representations take the form of shapelets that discriminate well one class from the others (Ye & Keogh, 2011). For instance, the Early Distinctive Shapelet Classification (Edsc) method learns a distance threshold for each shapelet, based on the computation of the Euclidean distance between the considered subsequence and all other valid subsequences in the training set (Xing et al., 2011). It selects a subset of them, based on a utility measure that combines precision and recall, weighted by the earliness. A prediction is made as soon as $\mathbf{x}_{t}$ matches one of these shapelets well enough. Because this family of methods is computationally expensive, extensions have been developed to reduce the computational load (Yan et al., 2020; Zhang & Wan, 2022). Other extensions aim at improving the reliability of the predictions (Ghalwash et al., 2014; Yao et al., 2019), and tackling multivariate time series (Ghalwash & Obradovic, 2012; He et al., 2013; 2015; Lin et al., 2015).
$\bullet$ The Earliest (Early and Adaptive Recurrent Label ESTimator) uses a RNN architecture (Hartvigsen et al., 2019) to make the prediction and a Reinforcement Learning agent trained jointly using policy gradient to trigger prediction or not. If a prediction is triggered, the hidden representation given by the RNN is sent to a Discriminator, whose role is to predict a class, given this representation. The model has been adapted to deal with irregularly sampled time series (Hartvigsen et al., 2022).
$\bullet$ Martinez et al. (2018; 2020) use a Deep Q-Network (Mnih et al., 2015), alongside a specifically designed reward signal, encouraging the agent to find a good trade-off between earliness and accuracy. Those types of approaches also naturally extend to online settings where time series are not of fixed length.
$\bullet$ The Decouple ETSC Network (Detscnet) (Chen et al., 2022) architecture leverages a gradient projection technique in order to jointly learn two sub-modules: one for variable-length series classification, and the other for the early exiting task.
$\bullet$ The Benefitter algorithm (Shekhar et al., 2023) is an anticipation-based approach that learns to predict the benefit of triggering a prediction early. This quantity is equal to the saving one could make by triggering some decision now minus the cost induced by a wrong prediction. A LSTM model is learned to regress the benefit, which thus triggers, at inference time, as soon as the benefit is positive, i.e. when savings induced by temporal costs exceed estimated misclassification costs.
$\bullet$ The Classifier-Induced Stopping (CIS) (Cao et al., 2023) model leverages policy gradient Reinforcement Learning in order to directly predict the optimal stopping time in a supervised way, found a posteriori for the training time series.
$\bullet$ The End-to-end Learned Early Classification of Time Series method (Elects) leverages an LSTM architecture, adding a stopping prediction head to the network and adapting the loss function to promote good early predictions (Rußwurm et al., 2023).
$\bullet$ Wang et al. (2024) introduce EarlyStop-RL, in which model-free RL is used to address the problem of early diagnosis of lung cancer.
C.2 End-to-end experiments
In this section, we test out two popular ECTS end-to-end methods, i.e. EARLIEST (Hartvigsen et al., 2019) and ELECTS (Rußwurm et al., 2023), and directly compare them to the separable baseline Proba Threshold. Due to the fact that those method do not directly expand to generic cost settings, we only perform this experiment using the standard cost setting, as described in Section 4.2.
Figure 9: Standard cost setting, non z-normalized proposed datasets (orange cylinder). It is found that end-to-end approaches outperform the separable baseline Proba Threshold for high temporal costs. Since end-to-end approaches are not constrained to consider time series only over the 20 timestamps corresponding to every 5% of the time series duration for prediction, they can make their prediction in the interval. For instance, before 5% of the time series has been observed, which happens for high delay cost, whereas separable methods, in this experimental setting, must wait until the first time step, resulting in higher average costs. However, as the delay cost decreases, end-to-end methods tend to be less stable overall, falling behind simple separable base method, such as Proba Threshold.
Appendix D Library
More details are given about the proposed library here. It leverages the modularity of separable ECTS approaches and offers the possibility to easily assess each of the component contribution to final performance. The main library object is an EarlyClassifier object that needs, at least, 3 inputs:
  • a ChronologicalClassifier that outputs classification prediction based on varying-size time series,
  • a TriggerModel that decides to whether accept or not the current classifier’s prediction,
  • a CostMatrices that defines the cost functions to be optimized.
As used throughout the paper, one way to implement a classification strategy is to use an ensemble of classifiers, each of which is specialized for a particular timestamp. The ClassifiersCollection object offers an interface to do that. It takes as input any scikit-learn estimator as well as a set of timestamps for which to learn a classifier. For example, the following code uses the same classification module used to produce Figure 8.
from sklearn.linear_model import RidgeClassifierCV
from sklearn.preprocessing import SplineTransformer
from sklearn.calibration import CalibratedClassifierCV
from sklearn.pipeline import make_pipeline
from ml_edm.classification.classifiers_collection import ClassifiersCollection
T # time series length
valid_timestamps = list(range(1, T+1)
clf = make_pipeline(
SplineTransformer(),
CalibratedClassifierCV(RidgeClassifierCV(), method="sigmoid")
collection_clf = ClassifiersCollection(
base_classifier=clf,
timestamps=valid_timestamps)
Then, having defined a classification strategy, one can fit a full ECTS model, for example, using the ProbabilityThreshold trigger model. By default, when no other argument than timestamp is given to the CostMatrices object, it uses the default cost setting, as defined in Section 4.2.
from ml_edm.early_classifier import EarlyClassifier
from ml_edm.trigger import ProbabilityThreshold
from ml_edm.cost_matrices import CostMatrices
early_clf = EarlyClassifier(
chronological_classifiers=collection_clf,
trigger_model=ProbabilityThreshold(valid_timestamps),
cost_matrices=CostMatrices(valid_timestamps),
early_clf.fit(X, y)
Thus, the library allows to easily pursue research in the separable ECTS field, facilitating the implementation of new approaches (whether in the classification or trigger part) and the systematic comparison with previous ones. It also pave the way for end-to-end approaches that can be framed as a trigger model that does not rely on any classifier’s outputs.
Appendix E Data description
E.1 UCR Time Series Classification datasets